A fertilizer advertised as 8-8-8 generally means it contains 8% total nitrogen, 8% P2O5, and 8% K2O.

Calculate the percentage of P and K in the fertilizer


1
Expert's answer
2021-04-19T07:31:43-0400

Solution:

Since the scale for percentages is 100, it is most convenient to calculate the mass of components present in a sample weighing 100 g. Taking this into account, the mass percentages provided may be more conveniently expressed as fractions:

8%N = 8 g N / 100 g sample

8% P2O5 = 8 g P2O5 / 100 g sample

8% K2O = 8 g K2O / 100 g sample


P2O5 → 2P

The molar mass of P2O5 is 141.9445 g/mol.

The molar mass of P is 30.974 g/mol.

%P' in P2O5 is:

%P' = (2 × 30.974 / 141.9445) × 100% = 43.64%

Mass of P = Mass of P2O5 × %P' / 100% = 8 g × 43.64% / 100% = 3.4914 g

%P in the fertilizer is:

%P = (Mass of P / Mass of sample) × 100%

%P = (3.4914 g / 100 g) × 100% = 3.4914% = 3.49%

%P in the fertilizer is 3.49%


K2O → 2K

The molar mass of K2O is 94.2 g/mol.

The molar mass of K is 39.0983 g/mol.

%K' in K2O is:

%K' = (2 × 39.0983 / 94.2) × 100% = 83.01%

Mass of K = Mass of K2O × %K' / 100% = 8 g × 83.01% / 100% = 6.6409 g

%K in the fertilizer is:

%K = (Mass of K / Mass of sample) × 100%

%K = (6.6409 g / 100 g) × 100% = 6.6409% = 6.64%

%K in the fertilizer is 6.64%


Answer:

%P in the fertilizer is 3.49%;

%K in the fertilizer is 6.64%.